OpenStudy (anonymous):

Any oscillating function can be written as \[ f(t)=k+\Re(\sum_{n=0}^{\infty}\check{c}_ne^{inwt} ) \] One can find t by \[ \check{c}_n=\frac{2\int_{0}^{T}f(t)e^{-inwt}}{T} \] I understood finding real Fourier coefficients for the real, but here why is there a 2 in the denominator?

OpenStudy (anonymous):

\[\check{c}_n=a_n-ib_n\] Where a is the cosine coefficient and b is the sine coefficient for the real fourier series, to clarify.

OpenStudy (experimentx):

can you give me the link were you got this?

OpenStudy (anonymous):

Feynman lectures

OpenStudy (anonymous):

For example, imagine a function\[f(t)=3\cos(6t)+12sint(6t)=\Re((3-12i)e^{i6t})\] We want to find 3-12i. \[\check{c}_n=\frac{2\int\limits_{0}^{2\pi}(3-12i)1dt}{2\pi}=\frac{2(3-12i)(2\pi)}{2\pi}=2(3-12i)\]

OpenStudy (experimentx):

it's pretty weird eq 26 http://mathworld.wolfram.com/FourierSeries.html

OpenStudy (anonymous):

So it seems the 2 was incorrect.

OpenStudy (experimentx):

\[ f(t)=3\cos(6t)+12\sin (6t)= 3 \; \Re (e^{i6t}) + 12 \Im ( e^{i6t}) \]

OpenStudy (anonymous):

Left hand side of that equation is only real- it can't have an imaginary part. Did you read a nonexistent i?

OpenStudy (experimentx):

Seems you eluded Imaginary part by putting complex coefficient.

OpenStudy (experimentx):

it would work out too.

OpenStudy (anonymous):

Sorry, I forgot taking the imaginary part takes the size of the imaginary part only without adding an I afterwards.

OpenStudy (experimentx):

3 - 12i <-- this is your coefficient. Your function is \( 3cos(6t)+12\sin(6t) \) try getting this out of that relation. I haven't tried this yet though.

OpenStudy (anonymous):

I tried it above, was I incorrect?

OpenStudy (experimentx):

your case, n=6 ... putting this on try this ... I'm not sure though \[ {3 \over \pi }\int_0^{\pi \over 3} (3\cos(6t)+12\sin(6t))e^{-i6t}dt = ?\]

OpenStudy (experimentx):

If that value = (3−12i) then your claim is correct.

OpenStudy (anonymous):

It should work also with 2pi also, which would make it a little easier. It is, just convert f(t) into ke^pt, \[(3-12i)e^{6t}\], as above, and the exponentials multiply to 1.

OpenStudy (anonymous):

I assume that if you 'complexify' a real function, operate on it, and take the real part of the result, it's the same as not complexifying it in the first place and operating just on reals for real result.

OpenStudy (experimentx):

\[ {3 \over \pi }\int_0^{\pi \over 3} (3\cos(6t)+12\sin(6t))e^{-i12t}dt =\]

OpenStudy (anonymous):

\[(3-12i)e^{i6t}e^{-i6t}=(3-12i)e^0=(3-12i)\]Am I missing something?

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=Integrate [%283+Cos[6+t]+%2B+12+Sin[6+t]%29%2FE^%28%286+I%29+t%29%2C+{t%2C+0%2C+Pi%2F3}]*3%2Fpi

OpenStudy (experimentx):

2 seems to be right!! or somehow i must get some other 2

OpenStudy (anonymous):

I don't see how...

OpenStudy (experimentx):

go to that wolf link.

OpenStudy (anonymous):

I saw that. Is my complexfying of 3cos(6t)+12sin(6t) right?

OpenStudy (experimentx):

that should be your coefficinet according to your formula. if you multiply by that by 2 ... you exactly get your coefficient and hence your claim is correct.

OpenStudy (experimentx):

since your function is very nice, you easily got this form \[ f(t)=3\cos(6t)+12sint(6t)=\Re((3-12i)e^{i6t}) ---- (1)\] if your function were not nice, you should have gone through more rigorous processes which is to do this \[ \check{c}_n=\frac{2\int_{0}^{T}f(t)e^{-inwt}}{T} ------ (2)\] the coefficient from (1) and (2) must match each other.

OpenStudy (experimentx):

and f(t) is not 3 - 12 i

OpenStudy (anonymous):

I wasn't proposing that. Why should the 2 exist?

OpenStudy (experimentx):

It's petty weird!!

OpenStudy (experimentx):

do you have mathematical derivation?

OpenStudy (anonymous):

Here's my logic why it shouldn't:\[F(t)=(3-12i)e^{i6t}\]\[\large \frac{2\int\limits_{0}^{\pi/3}(3-12i)e^{i6t}e^{-i6t}dt}{\pi/3}= \frac{2\int\limits_{0}^{\pi/3}(3-12i)dt}{\pi/3}=\frac{2(3-12i)\int\limits_{0}^{\pi/3}dt}{\pi/3}\]\[=\frac{2(3-12i)\pi/3}{\pi/3}=2(3-12i)\]

OpenStudy (experimentx):

All fourier coefficients have while not expressing into complex form. May be due to the fact that you are using Re(....) there.

OpenStudy (experimentx):

you are on wrong track

OpenStudy (anonymous):

You may be correct.I think my logic rests on this, is it incorrect:\[g(t)=\Re(h(t))\]\[F(g(t))=\Re(F(h(t))\]?

OpenStudy (experimentx):

http://www.wolframalpha.com/input/?i=Re [%283-12i%29e^%28i6t%29] these two are just two different form of same equation.

OpenStudy (experimentx):

you already have done this stuff ... the only thing you need to prove is